how do I solve
2x^3+3x^2-65x+84+0

Answer by stanbon(75887)

solve
2x^3+3x^2-65x+84 = 0
----
I graphed it and found a rational zero at x = -7
---
Then, using synthetic division, I got:
-7)....2....3....-65....84
.......2.....-11...12...|..0
----
Quotient: 2x^2-11x+12 = 0
2x^2-8x-3x+12 = 0
2x(x-4)-3(x-4) = 0
(x-4)(2x-3) = 0
x = 4 or x = 3/2
----
Final answer: x = -7 or x = 3/2 or x = 4
============================================
Cheers,
Stan H.